OpenStudy (anonymous):

Hi! i want increase ampere at AC current without decrease voltage . so what element can do this? e.g: can armature winder do this? thank you for your help

OpenStudy (anonymous):

Connect a capacitor in series with a resistor in your circuit, similar to an RC circuit. For example, an RC circuit consists of a voltage supply, "Vs," connected in series with a resistor, "R," and a capacitor, "C," connected in series with the resistor. Calculate the current using the formula I (t) = Vs/R * e^-t/RC where t stands for the elapsed time since the power supply, Vs, was turned on and I(t) means current that changes with the elapsed time. For example, if Vs is 120 volts, R is 300 ohms, C is 5 nanofarads and t = 3 microseconds: C = 5 nanofarads or 0.000000005 farads or 5 x 10^-9 RC = (300)(5 x 10^-9) = 0.0000015 or 1.5 microseconds. Convert to microseconds to get RC in the same units as t. I (t) = 120/300 * e^-3/1.5= 0.4 (e^-2) = (0.4)(0.8) = 0.32 amps. Increase the current level by increasing the value of the capacitor. For example, increase the value of the capacitor from 5 nanofarads to 5 microfarads: C = 5 microfarads = 0.000015 farads = 5 x 10^-6. RC = (300)(5 x 10^-6) = 0.0015 or 1,500 microseconds. I (t) = 120/300 * e^-3/1500 = 0.4 (e^-0.002) = (0.4)(0.998) = 0.3992. The current increases as you increase the value of the capacitor.